# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v4

@@ Appendix B Proofs for Results in Main Text
Conversely, note that for any $\tau>0$ and any $\varepsilon>0,$ 5.2 implies that there exist $\kappa$ and (random) $\Omega_{+},\Omega_{-}\in\mathcal{W}_{\kappa}$ such that $\hat{\sigma}_{\Omega_{+}},$ $\hat{\sigma}_{\Omega_{-}}\leq\sqrt{1+\tau^{2}}\hat{\sigma}_{\Sigma^{-1}}$ and
$\limsup_{n\to\infty}\operatorname{P}\left\{\sqrt{n}\left|\hat{\theta}_{\Omega_{+}}-\hat{\theta}_{\Sigma^{-1}}-\sqrt{J}\hat{\sigma}_{\Sigma^{-1}}\tau\right|>\varepsilon\right\}<\varepsilon$
$\limsup_{n\to\infty}\operatorname{P}\left\{\sqrt{n}\left|\hat{\theta}_{\Omega_{-}}-\hat{\theta}_{\Sigma^{-1}}+\sqrt{J}\hat{\sigma}_{\Sigma^{-1}}\tau\right|>\varepsilon\right\}<\varepsilon$
@@
$\limsup_{n\to\infty}\operatorname{P}\left\{\left|t_{\Omega_{+}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon^{\prime}\right\}<\varepsilon^{\prime},$
$\limsup_{n\to\infty}\operatorname{P}\left\{\left|t_{\Omega_{-}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon^{\prime}\right\}<\varepsilon^{\prime}.$
@@
However, since $|t_{\Omega_{+}}(\theta_{0})|$ is increasing in $\theta_{0}$ for $\theta_{0}\leq\hat{\theta}_{\Sigma^{-1}}$ while $|t_{\Omega_{+}}(\theta_{0})|$ is increasing in $\theta_{0}$ for $\theta_{0}\geq\theta_{\Sigma^{-1}},$
@@
However, we have shows that for $\kappa$ sufficiently large, we can lower bound the right hand side by $\sqrt{J}-\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0.$ Thus, for all $\varepsilon>0$ and $\tau>0,$
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left\{\inf_{\theta_{0}}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}(\theta_{0})\right|<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon\right\}=0.$
@@
The first part of the result is immediate from Corollary 5.3. For the second part, note that the proof of Corollary 5.3 shows that $\left|t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)\right|\leq\sqrt{J^{*}}$ for all $\Omega,$ while by (9), $t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)-t_{\Omega}(\hat{\theta}_{\Sigma^{-1}})\overset{p}{\longrightarrow}0$ uniformly over $\Omega\in\mathcal{W}_{\kappa}$ and $\sqrt{J^{*}}-\sqrt{J}\overset{p}{\longrightarrow}0.$ Thus,
$\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}\right)-\sqrt{J}\right|\overset{p}{\longrightarrow}0,$
@@
On the other hand, note that since $C_{\Omega}(c)$ is an interval for all $c,$ $\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(\sqrt{J}+\nu)$ is also an interval. As argued in the proof of Corollary 5.3, for any $\tau>0$ and any $\nu^{\prime}>0$ there exists $\kappa>0$ and (random) $\Omega_{+}$ and $\Omega_{-}\in\mathcal{W}_{\kappa}$ such that $\hat{\sigma}_{\Omega_{+}},$ $\hat{\sigma}_{\Omega_{-}}\leq\sqrt{1+\tau^{2}}\hat{\sigma}_{\Sigma^{-1}}$ and
$\limsup_{n\to\infty}\operatorname{P}\left\{\left|t_{\Omega_{+}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\nu^{\prime}\right\}<\nu^{\prime},$
$\limsup_{n\to\infty}\operatorname{P}\left\{\left|t_{\Omega_{-}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\nu^{\prime}\right\}<\nu^{\prime}.$
As also argued in the proof of Corollary 5.3, for all $\theta_{0}$,
@@
$\limsup_{n\to\infty}\operatorname{P}\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}<2\nu^{\prime}.$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}=1.$
@@
$\lim_{\nu\to 0}\sup_{\tau>0}\liminf_{n\to\infty}\operatorname{P}\left\{\sqrt{J}+\nu<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}\right\}=1,$
@@
$\lim_{\nu\to 0}\lim_{\kappa\to\infty}\limsup_{n\to\infty}\operatorname{P}\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\sqrt{J}+\nu\right\}=1.$

### v5

@@ Appendix B Proofs for Results in Main Text
Conversely, note that for any $\tau>0$ and any $\varepsilon>0,$ 5.2 implies that there exist $\kappa$ and (random) weight matrices $\Omega_{+},\Omega_{-}\in\mathcal{W}_{\kappa}$ such that $\hat{\sigma}_{\Omega_{+}},$ $\hat{\sigma}_{\Omega_{-}}\leq\sqrt{1+\tau^{2}}\hat{\sigma}_{\Sigma^{-1}}$ and $\Omega_{+},\Omega_{-}$ approximately obtain the bounds in the proposition,
(13) $\begin{split}&\limsup_{n\to\infty}P\left\{\sqrt{n}\left|\hat{\theta}_{\Omega_{+}}-\hat{\theta}_{\Sigma^{-1}}-\sqrt{J}\hat{\sigma}_{\Sigma^{-1}}\tau\right|>\varepsilon\right\}<\varepsilon,\\ &\limsup_{n\to\infty}P\left\{\sqrt{n}\left|\hat{\theta}_{\Omega_{-}}-\hat{\theta}_{\Sigma^{-1}}+\sqrt{J}\hat{\sigma}_{\Sigma^{-1}}\tau\right|>\varepsilon\right\}<\varepsilon.\end{split}$
@@
(14) $\begin{split}&\limsup_{n\to\infty}P\left\{\left|t_{\Omega_{+}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon^{\prime}\right\}<\varepsilon^{\prime},\\ &\limsup_{n\to\infty}P\left\{\left|t_{\Omega_{-}}(\hat{\theta}_{\Sigma^{-1}})\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon^{\prime}\right\}<\varepsilon^{\prime}.\end{split}$
@@
However, since (with high probability) $|t_{\Omega_{+}}(\theta_{0})|$ is decreasing in $\theta_{0}$ for $\theta_{0}\leq\hat{\theta}_{\Sigma^{-1}}$ while $|t_{\Omega_{-}}(\theta_{0})|$ is increasing in $\theta_{0}$ for $\theta_{0}\geq\theta_{\Sigma^{-1}},$
@@
However, we have shown that for $\kappa$ sufficiently large, with high probability we can lower bound the right hand side by $\sqrt{J}-\varepsilon^{\prime}$ for any $\varepsilon^{\prime}>0.$ Thus, for all $\varepsilon>0$ and $\tau>0,$
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left\{\inf_{\theta_{0}}\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}(\theta_{0})\right|<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}-\varepsilon\right\}=0.$
@@
The first part of the result is immediate from Corollary 5.3. For the second part, note that the proof of Corollary 5.3 shows that $\left|t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)\right|\leq\sqrt{J^{*}}$ for all $\Omega,$ while by (9), $t^{*}_{\Omega}\left(\vartheta(\psi)+\frac{1}{\sqrt{n}}\hat{\theta}^{*}_{\Sigma^{-1}}\right)-t_{\Omega}(\hat{\theta}_{\Sigma^{-1}})\overset{p}{\longrightarrow}0$ uniformly over $\Omega\in\mathcal{W}_{\kappa}$ and $\sqrt{J^{*}}-\sqrt{J}\overset{p}{\longrightarrow}0.$ Thus, for all $\varepsilon^{\prime}$
$\limsup_{n\to\infty}P\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}\right)-\sqrt{J}\right|>\varepsilon^{\prime}\right\}\to 0,$
@@
On the other hand, note that since $C_{\Omega}(c)$ is an interval for all $c,$ $\bigcap_{\Omega\in\mathcal{W}_{\kappa}}C_{\Omega}(\sqrt{J}+\nu)$ is also an interval. For any $\tau>0$ and any $\nu^{\prime}>0$, following the proof of Corollary 5.3, there exists $\kappa\geq 1$ and (random) $\Omega_{+}$ and $\Omega_{-}\in\mathcal{W}_{\kappa}$ such that $\hat{\sigma}_{\Omega_{+}},$ $\hat{\sigma}_{\Omega_{-}}\leq\sqrt{1+\tau^{2}}\hat{\sigma}_{\Sigma^{-1}}$, and (LABEL:eq:omega+-def) and (LABEL:eq:tstat+-bound) hold (with $\nu^{\prime}$ in place of $\varepsilon$ and $\varepsilon^{\prime}$). As also argued in the proof of Corollary 5.3, for all $\theta_{0}$,
@@
$\limsup_{n\to\infty}P\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|\leq\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}<2\nu^{\prime}.$
@@
$\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}-\nu^{\prime}\right\}=1.$
@@
$\lim_{\nu\to 0}\sup_{\tau>0}\liminf_{n\to\infty}P\left\{\sqrt{J}+\nu<\frac{\tau}{\sqrt{1+\tau^{2}}}\sqrt{J}+\frac{|\delta|}{\sqrt{1+\tau^{2}}\sigma_{\Sigma^{-1}}}\right\}=1,$
@@
$\lim_{\nu\to 0}\lim_{\kappa\to\infty}\limsup_{n\to\infty}P\left\{\sup_{\Omega\in\mathcal{W}_{\kappa}}\left|t_{\Omega}\left(\hat{\theta}_{\Sigma^{-1}}+\frac{1}{\sqrt{n}}\delta\right)\right|>\sqrt{J}+\nu\right\}=1.$